Main INDEX
Monthly INDEX
PREV
NEXT
User name H. Ibrahim
Log entry time 00:05:33 on November14,2002
Entry number 90062
keyword=Swing Shift Summary
Swing shift crew: K. McCormick and H. Ibrahim
Shift Summary
----------------
16:50 MCC called to inform us that hall C will get beam soon and they have to take the beam from us until they do that.
17:00 Beam is gone.
17:15 Beam is back.
17:45 MCC said that they see many beam trips and they will take the beam to check the problem. No beam right now.
17:52 Hall C to controlled access.
17:58 Called MCC to know the beam status and they said that there is a radiation monitor problem in the transport line to hall C which is causing this. No time estimate at this moment. Stoped the current run (#3151).
19:00 Called MCC to know the current situation and they said that they need another 15 min before they give us the beam.
19:35 MCC called to see if we are ready to recieve beam.
19:45 Beam is back at 100 uA.
20:30 MCC said that hall C is still causing beam trips. Beam is gone for now.
20:50 Called MCC to get an estimate of the time they need to give us the beam back and they told us that there is beam instability in the injector and they have called an expert to invistigate the problem. Nothing is expeted until then. Stoped current run (#3152).
21:25 Beam is back at 100 uA.
22:45 We noticed some run control error messages with many beeps and zero data rate. We stoped the current run (#3154) and began to check the source of the problem and while we were doing that everything (amazingly) became fine without any need to reset the ROCs. We started a new run and everything looks good.
24:00 End of the swing shift.
Run Summary
----------------
Continuing Q3D_j40 kinematics:
Run # 3150 (Production, 1 M events)
Run # 3151 (Production, 487 k events)
Run # 3152 (Production, 738 k events)
Run # 3153 (Production, 1 M events)
Run # 3154 (Production, 438 k events)
Run # 3155 (Production, 1 M events)
Run # 3156 (Production, in progress)